OpenStudy (clamin):

PLEASE HELP!! MEDAL!! HOW COME THE ANSWER FOR 3n+1=2 IS 1/3??

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

subtract 1, divide by 3, that is what you get

OpenStudy (dtan5457):

3(1/3)+1=2 1/3 x 3=1 1+1=2

OpenStudy (clamin):

oh okay now i get it @misty1212

OpenStudy (ny,ny):

3n + 1 = 2 - 1 -1 ---------- 3n = 1 -- -- 3 3 n = 1/3 to solve for n, you have to isolate everything besides it. subtract 1 on both sides. divide by 3 on both sides. equals 1/3.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (misty1212):

if i can make it there...